OpenStudy (anonymous):

A line contains (9, y) and (-6, 3). It has a slope of 2/3. find y. PLEASE SHOW STEPS AND EXPLAIN REASONING!!! :)

myininaya (myininaya):

\[\frac{3-y}{-6-9}=\frac{2}{3}\]

OpenStudy (anonymous):

y = mx + b y = 2/3x + b 3 = (2/3)(-6) + b b = 7 y =(2/3)x + 7 y = 13

myininaya (myininaya):

-6-9=-15 multiply -15 on both sides then subtract 3 on both sides then multiply -1 on both sides

OpenStudy (anonymous):

myininaya what is the answer for y of your first response?

myininaya (myininaya):

follow these steps to solve for y if you have trouble on a step let me know also i will be happy to check your work

OpenStudy (anonymous):

are you a teacher?

myininaya (myininaya):

why do you ask?

myininaya (myininaya):

cow has a cute way of solving for y too go cow!